# Symbolic Learning of Interpretable Reduced-Order Models for Jumping Quadruped Robots

### v1

@@ 3.3 Data Collection and Preprocessing
Datasets were collected from both simulated and real-world environment. Simulated data were generated using the PyBullet engine (Coumans and Bai, 2016) with Unitree Go1 and A1 robots. Real-world data were gathered from an actual Go1 robot to validate the simulation results, ensure practical applicability, and benchmark the algorithm’s performance on real-world, noisy data.
Simulation data for the Go1 quadruped robot captured forward jumping motions, as shown in Fig. 4(a), achieved with Model Predictive Control (MPC) (Ding et al., 2024a). Additionally, data from the Unitree A1 quadruped for froggy jumping behavior, as visualized in Fig. 4(b), were collected using a Cartesian PD controller (Nguyen and Nguyen, 2022). For both robots, the dataset includes joint angles and velocities ($\mathbf{q}_{\mathrm{j}},\dot{\mathbf{q}}_{\mathrm{j}}\in\mathbb{R}^{12}$), body position and velocity ($\mathbf{q}_{\mathrm{b}},\dot{\mathbf{q}}_{\mathrm{b}}\in\mathbb{R}^{6}$), input torques ($\bm{\tau}\in\mathbb{R}^{12}$), and the contact state ($\mathbf{c}\in\mathbb{R}^{4}$) for each timestep.
The Go1 simulated data consists of 156 jumps (120 for training, 30 for validation, and 6 for testing), with each jump having 800 data points: 300 for launching, 300 for landing (totaling 600 contact points), and 200 for flight. The A1 simulated dataset contains 150 jumps (100 for training, 40 for validation, and 10 for testing), with each jump comprising 1200 data points: 400 for launch, 400 for landing (800 contact points), and 400 for flight. To enhance robustness, a variant of the datasets with added noise as described in the work from Ding et al. (2024a).
Figure 4: A sequential images of quadruped jumps: (a) A simulated Go1 robot performs a vertically-oriented jump. (b) A simulated A1 executes a froggy jump. (c) A Go1 robot performs a (vertically-oriented) jump in the real world.
Real-world data were collected from a Go1 quadruped, as depicted in Fig. 4(c). The quadruped was operated using the same control algorithm as in the simulation, with data collection performed similarly using the official Go1 SDK. This dataset comprises 15 jumps, distributed as 12 for training, 2 for validation, and 1 for testing, maintaining the same data point distribution as the Go1 simulated data.
@@ 3.4 Evaluation Procedure
In this study, we regard the aSLIP model (see Appendix.A) as the baseline for predicting the dynamical behavior.
4 Results
This section evaluates the approach using simulated Unitree Go1 and A1 quadrupeds data and real-world data from a Unitree Go1.
4.1 Go1 Simulation Results
4.1.1 Learning in 2D Latent Space for Go1
We initially explored a two-dimensional latent model for the synchronized jumping. Fig. 5 visualizes the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the trained encoder, indicating that the (learned) latent dimensions primarily correspond to the robot’s body center positions in the $x$ and $z$ coordinates.
Figure 5: Heatmap of the encoder’s transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ trained on the simulated Go1 dataset with two latent dimensions. Brighter colors indicate higher absolute weights. The vertical axis shows the 18 configuration dimensions, and the horizontal axis shows the two latent dimensions.
$\begin{cases}\ddot{\xi}_{1}\!\!=\!\!0.36\!-\!0.16\,\dot{\xi}_{1}\!-\!\!0.91\,\dot{\xi}_{2}\!-\!0.75\sin(\dot{\xi}_{1})\!+\!0.11\,\nu_{1}\!-\!0.05\,\nu_{2}\\ \ddot{\xi}_{2}\!\!=\!\!-0.16\!-\!0.23\,\dot{\xi}_{1}\!-\!0.75\,\dot{\xi}_{2}\!-\!0.96\,\sin(\dot{\xi}_{2})\!+\!0.14\,\nu_{2}\end{cases},$ (12)
and
$\begin{cases}\ddot{\xi}_{1}=0.29-1.22\,\dot{\xi}_{1}-1.00\sin(\dot{\xi}_{1})+51.05\,\nu_{1}\\ \ddot{\xi}_{2}=0.42\,\dot{\xi}_{1}+0.52\sin(\dot{\xi}_{1})\end{cases}.$ (13)
The prediction results for simulated Go 1 jumping are shown in Fig. 6.
Figure 6: Evaluation of the trained model with two latent dimensions on the simulated Go1 jumping dataset.
Fig. 7 compares the body’s linear position along the $x$ and $z$ axes, demonstrating that the learned model outperforms, especially in the $x$ axis, the baseline aSLIP model.
Figure 7: Comparison between the learned model and the aSLIP model, on the task of predicting the CoM linear positions along the $x$ and $z$ axes, where the simulated Go 1 performs a jumping motion.
@@ 4.1.2 Latent Space Dimensionality Tradeoff Analysis
We analyze the effect of latent dimensionality on model design. As shown in Fig. 8, increasing to 18 dimensions does not change the dominance of key full-order components.
Figure 8: Heatmap visualization of the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the encoder trained on the simulated Go1 dataset in the case of 181818 latent dimensions.
Fig.9 further evaluates the trade-off between accuracy (via final reconstruction error $\mathcal{E}_{\text{dec}}$) and complexity (via active SINDy coefficients $|\bm{\Xi}|$) across 18 contact phase models with varying latent dimensions.
Figure 9: The trade-off between accuracy and complexity with increasing latent dimensions evaluated on the simulated Go1 dataset. The error bars depict the standard deviation from five different random seeds.
To quantitatively evaluate this trade-off, we employ a combined loss $\mathcal{L}_{\text{mod}}$ that is inspired by the Akaike Information Criterion (AIC) for model selection (Bozdogan, 1987) and calculated as
@@
with $\lambda=0.001$, balancing the loss terms, the combined loss (red line in Fig.9) suggests optimal performance with 4–6 latent dimensions. This aligns with the visibly significant components in Fig.8, where further increasing dimensionality yields marginal gains at the cost of complexity.
@@ 4.2 A1 Simulation Results
We also conducted experiments on data collected with a simulated Unitree A1 quadruped executing a froggy jump to evaluate the algorithm’s effectiveness across different platforms. The results are illustrated in Fig. 10.
Figure 10: Evaluation of the proposed method with four latent dimensions on A1 froggy jumping.
A notable difference is observed in the body angular position $y$ between the Unitree A1 and the Unitree Go1. The froggy jump causes a more pronounced angular rotation along the $y$-axis in the A1’s case. Despite the increased significance of this angular dimension for the jump motion, the four-dimensional model successfully captures this dynamic and most joint positions.
@@ 4.3 Go1 Real World Results
Given the scarcity of real-world compared to simulated data, we first train the model on simulation data with added noise and then subsequently fine-tune it on the few-sample real-world dataset. Fig. 11 illustrates the model’s ability to replicate a jump on the hardware, comparing joint angles and CoM positions and orientations.
Figure 11: Evaluation of the learned reduced-order model with four latent dimensions on the Go1 real-world testing dataset.
The proposed model performs well on actual hardware data, despite reduced accuracy during flight and initial landing phases compared to simulations. Nonetheless, the model effectively manages real-world noise and complexity, demonstrating robust performance.
@@ 5 Discussion and Conclusion
The results of our symbolic learning framework show that a reduced latent space can effectively preserve key dynamic patterns of the quadruped robot jumping motion. By combining autoencoder-based embeddings with sparse regression, we obtain interpretable and compact models that approximate the complex full-order dynamics. Our analysis of latent dimensionality shows that increasing beyond 4–6 dimensions offers diminishing returns in accuracy while introducing additional model complexity. The sim-to-real transfer results indicate that the model retains core dynamic features and generalizes well despite limited real-world data, validating its robustness to measurement noise and domain shift.
In future work, we aim to extend this methodology to more dynamic behaviors, such as walking and running, and to systems involving rapid contact changes or higher-dimensional morphologies, including humanoids.

### v2

@@ 3.3 Data Collection and Preprocessing
Datasets were collected from both simulated and real-world environments. Simulated data were generated using the PyBullet engine (Coumans and Bai, 2016) with Unitree Go1 and A1 robots. Real-world data were gathered from an actual Go1 robot.
Simulation data for Go1 captured forward pronking motions (see Fig. 4(a)), achieved with model predictive control (MPC) (Ding et al., 2024a). Data from A1 for froggy jumping (see Fig. 4(b)) were collected using a Cartesian PD controller (Nguyen and Nguyen, 2022). The dataset includes joint angles and velocities ($\mathbf{q}_{\mathrm{j}},\dot{\mathbf{q}}_{\mathrm{j}}\in\mathbb{R}^{12}$), body position and velocity ($\mathbf{q}_{\mathrm{b}},\dot{\mathbf{q}}_{\mathrm{b}}\in\mathbb{R}^{6}$), input torques ($\bm{\tau}\in\mathbb{R}^{12}$), and the contact state ($\mathbf{c}\in\mathbb{R}^{4}$) for each timestep.
The Go1 simulated data consists of 156 jumps (120 for training, 30 for validation, and 6 for testing), with each jump having 800 points: 300 for launching, 300 for landing, and 200 for flight. The A1 dataset contains 150 jumps (100 for training, 40 for validation, and 10 for testing), with each jump comprising 1200 points: 400 for launch, 400 for landing, and 400 for flight. To enhance robustness, a variant of the datasets with added noise was considered, as described in the work from Panichi et al. (2025).
Figure 4: Sequential images of quadruped jumps: (a) A simulated Go1 robot performs a pronking jump. (b) A simulated A1 executes a froggy jump. (c) A real Go1 performs a pronking jump.
Real-world data were collected from a Go1 robot, as depicted in Fig. 4(c). The quadruped was operated using the same control algorithm as in the simulation, with data collection performed similarly using the official Go1 SDK. This dataset comprises 15 jumps, distributed as 12 for training, 2 for validation, and 1 for testing, maintaining the same data point distribution as the Go1 simulated data.
@@ 3.4 Evaluation Procedure
To validate the model $\Theta$, the initial state $(\mathbf{q}_{0},\dot{\mathbf{q}}_{0})$ and inputs $\mathbf{u}$ are transformed by the trained encoder and the input transformation equation to the latent initial state $(\bm{\xi}_{0},\dot{\bm{\xi}}_{0})$ and $\bm{\nu}$, respectively. The model then predicts latent accelerations $\ddot{\bm{\xi}}_{\text{pred}}$, which are integrated using the lsoda method (Hindmarsh and Petzold, 2005) to obtain predicted latent space state ${\bm{\xi}}_{\text{pred}}$ and $\dot{\bm{\xi}}_{\text{pred}}$ with a step size of $500\text{\,}\mathrm{H}\mathrm{z}$. This process is repeated for each time step. The predicted latent configuration $\bm{\xi}_{\text{pred}}$ is then decoded into a original configuration space by $\mathbf{q}_{\text{pred}}=\mathbf{W}_{\mathrm{d}}\,\bm{\xi}_{\text{pred}}+\mathbf{B}_{\mathrm{d}}$ for direct comparison with actual jump data.
An alternative evaluation approach involves simulating the learned system for shorter intervals (e.g., $0.1\text{\,}\mathrm{s}$ – 50 steps) with periodic resets to the original latent state $(\bm{\xi},\dot{\bm{\xi}})$. This strategy decreases error accumulation over time, showcasing the precision of the model for medium-horizon forecasts, which could be used in planning and control frameworks such as MPC.
4 Results
This section evaluates the approach. We explored several latent-space dimensionalities to assess the trade-off between model compactness and accuracy. Two representative configurations are reported in Table 1.
Table 1: Hyperparameters of the models. The number of candidate functions is computed as $\sum_{k=0}^{p}(2l+k-1)+2(2l)+l$ for latent dimension $l$ and polynomial order pp.
Parameter Symbol / Value Go1 A1
Latent dimension $l$ 2 4
Polynomial order pp 1 1
# Candidates $|\Theta|$ 15 29
Sparsity threshold $\lambda$ 0.01
Learning rate $\eta$ $1\times 10^{-3}$
Batch size $B$ 100
Epochs $N_{\text{epoch}}$ 501
4.1 Go1 Simulation Results
4.1.1 Learning Go1 pronking
We initially explored a two-dimensional latent model for the pronking. Fig. 5 visualizes the transposed weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ of the trained encoder, indicating that the latent dimensions primarily correspond to the CoM positions in the $x$ and $z$ coordinates.
Figure 5: Heatmap of the encoder weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ (Go1 simulation, two latent dimensions). Brighter colors indicate larger weights; rows are 18 configuration dimensions and columns are the two latent dimensions.
$\begin{cases}\ddot{\xi}_{1}\!\!=\!\!0.36\!-\!0.16\,\dot{\xi}_{1}\!-\!\!0.91\,\dot{\xi}_{2}\!-\!0.75\sin(\dot{\xi}_{1})\!+\!0.11\,\nu_{1}\!-\!0.05\,\nu_{2},\\ \ddot{\xi}_{2}\!\!=\!\!-0.16\!-\!0.23\,\dot{\xi}_{1}\!-\!0.75\,\dot{\xi}_{2}\!-\!0.96\,\sin(\dot{\xi}_{2})\!+\!0.14\,\nu_{2},\end{cases}$ (12)
and
$\begin{cases}\ddot{\xi}_{1}=0.29-1.22\,\dot{\xi}_{1}-1.00\sin(\dot{\xi}_{1})+51.05\,\nu_{1},\\ \ddot{\xi}_{2}=0.42\,\dot{\xi}_{1}+0.52\sin(\dot{\xi}_{1}).\end{cases}$ (13)
The latent coordinates do not correspond to explicit physical variables, but the learned symbolic equations display structures that are consistent with key principles of quadruped jumping. The constant terms behave like gravity-related biases, the linear velocity terms act as effective damping, and the input-dependent components reflect how actuation influences the motion. The trigonometric functions of latent velocities serve as smooth nonlinear shaping terms that encode phase-dependent effects arising from posture changes, contact transitions, and limb coordination. These nonlinear components naturally emerge when compressing high-dimensional multibody dynamics into a low-dimensional latent representation.
Figure 6: Evaluation of the trained model (with two latent dimensions) for simulated Go1 pronking. Considering the symmetry between right and left side, we only report the joints angle profiles in front left (fl) and rear right ($r$) leg.
Table 2: Prediction performance of the trained model on simulated and real datasets. Values denote mean absolute error $\pm$ standard deviation for joint angles and CoM pose components.
Dataset Pre. Joint angles CoM position CoM orientation
fl hx fl hy fl hz rr hx rr hy rr hz x y z roll pitch yaw
@@ 4.1.1 Learning Go1 pronking
[…]
The prediction results for Go1 pronking are shown in Fig. 6. Table 2 reports the mean errors and the standard deviation (std) along each direction, demonstrating the high prediction accuracy. To furthermore demonstrate the effectiveness, we regard the actuated SLIP (aSLIP) model (see Appendix.A) as the baseline. Fig. 7 compares the body’s linear position along the $x$ and $z$ axes, demonstrating that the learned model outperforms the baseline reduced-order model. Statistical results demonstrate that the average prediction error is reduced from 5.7e-2 m to 1.6e-2 m along the $x$ direction and reduced from 9.5e-2 m to 4.0e-2 m along $z$ direction.
Figure 7: Comparison between the learned model and the aSLIP model regarding the CoM positions prediction, where the Go 1 performs a pronking jumping motion.
@@ 4.1.2 Latent Space Dimensionality Tradeoff Analysis
Fig. 8 illustrates the effect of latent dimensionality on model design.
Figure 8: Heatmap of the encoder weight matrix $\mathbf{W}_{\mathrm{e}}^{\mathrm{T}}$ (Go1 simulation, 18 latent dimensions).
Fig. 9 further evaluates the trade-off between accuracy (via final reconstruction error $\mathcal{E}_{\text{dec}}$) and complexity (via active coefficients $|\bm{\Xi}|$) across 18 latent dimensions.
Figure 9: Accuracy vs. complexity with the increase of latent space dimensions.
To quantitatively evaluate this trade-off, we employ a combined loss $\mathcal{L}_{\text{mod}}$ (Bozdogan, 1987), calculated as
@@
with $\lambda=0.001$.
The combined loss (red line in Fig. 9) suggests optimal performance with 4–6 latent dimensions. This aligns with the visibly significant components in Fig. 8, where further increasing dimensionality yields marginal gains at the cost of complexity.
@@ 4.2 A1 Simulation Results
We also conducted experiments with a simulated A1 froggy jump to evaluate the algorithm’s effectiveness across different platforms and modalities. The results are illustrated in Fig. 10 and Table 2.
Figure 10: Learning A1 froggy jumping (simulation data) with four latent dimensions.
A notable difference is observed in the body angular position $y$ (i.e., pitch angle) between the Unitree A1 and Go1. The froggy jump causes a more pronounced angular rotation along the $y$-axis in the A1’s case. Despite the increased significance of this angular dimension for the froggy jump, the four-dimensional model successfully captures this dynamic and most joint positions. High precision is obtained when reset the latent state every 0.1 s.
@@ 4.3 Go1 Real World Results
Given the scarcity of real-world compared to simulated data, we first train the model on simulation data with added noise and then subsequently fine-tune it on the few-sample real-world dataset. Fig. 11 illustrates the model’s ability to replicate a jump on the hardware, and Table 2 reports the prediction error.
Figure 11: Learning Go1 pronking jumping (real data) with four latent dimensions.
As can be seen from Fig. 11 and Table 2, the proposed model performs well on actual hardware data, despite reduced accuracy during flight and initial landing phases compared to simulations. The divergence between real and predicted trajectories mainly results from unmodeled actuator and contact effects that accumulate during real execution. Applying a domain randomization in the future work strategy further alleviates this discrepancy by enhancing the model’s robustness to such unmodeled variations. The symbolic model still captures the key motion patterns, with residual errors from friction, delay, and noise remaining bounded.
@@ 5 Discussion and Conclusion
This work demonstrates that a low-dimensional, physics-aligned latent space can faithfully capture the essential dynamics of quadruped jumping. By combining a linear autoencoder embedding with sparse regression, the proposed framework yields compact and interpretable symbolic models that approximate full-order dynamics with meaningful accuracy. The learned equations recover gravity-like loading, effective damping, and actuation influence, while revealing coupling effects absent in idealized template models yet necessary to represent the richer behaviors of the full quadruped system. Beyond these empirical results, this study highlights how a linear latent embedding—together with a coherent multi-phase training strategy—enables SINDy to discover reduced-order template dynamics for hybrid legged motion, marking a conceptual departure from prior nonlinear SINDy-AE frameworks.
The approach remains sensitive to unmodeled actuator and contact effects, which accumulate during long-horizon rollouts and partly explain the observed sim-to-real discrepancies. Future work will extend the methodology to more diverse locomotion behaviors, incorporate hybrid contact transitions more explicitly, and scale the framework to higher-dimensional morphologies. Another promising direction is to integrate the learned symbolic models into adaptive or phase-dependent controllers for real-time robot movement tasks.